Fermi National Laboratory
 

Accelerator Update for May 20, 2009 - May 22, 2009

 

Wednesday May 20

The day shift began with Operations monitoring the Tevatron (TeV) store 7066 (118.0E30) and the Recycler (33.3E10), with the Antiproton Source stacking (24.8E10), with the Main Injector (MI), Linac, and Booster operating normally, with the Meson MTest, MiniBooNE, and NuMI experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

Operations terminated store 7066 at 2 PM. D0 personnel accessed their collision hall following store termination.

Operations reported that the Main Injector quadrupole focusing bus had a ground fault. Investigation found the quadrupole magnet at the MI-428 location had developed a substantial LCW leak. An attempt was made to braze the copper joint, but the leak remained. The Run Coordinator decided to replace the magnet. Support personnel had located a replacement magnet by the end of the evening shift.

Thursday May 21

The midnight shift began with Operations monitoring the stack (12.1E10).

At the beginning of the midnight shift, Mechanical and EE Support personnel accessed the Main Injector and began the magnet replacement. The new magnet was in the tunnel by the end of shift.

Main Injector quadrupole magnet installation continued through out the day shift.

The TeV accesses concluded at 12:50 PM.

Operations put the TeV through a dry squeeze at 3:05 PM.

The MI magnet replacement was finished by 4:40 PM.

MI was ramping by 5:25 PM.

Operations established protons to Pbar and NuMI by 5:40 PM.

Operations established beam to MTest at 7:10 PM.

An air conditioning problem inhibited MiniBooNE beam.

TeV P\personnel began a study period at 8:30 PM.

Operations called in Booster RF personnel to investigate the failure of two RF stations.

Friday May 22

The midnight shift began with Operations monitoring the stash (63.7E10), with Pbar stacking (26.4E10), and with NuMI taking beam. MiniBooNE was off due to a HVAC problem.

TeV studies concluded at 3:05 AM.

Operations put the TeV into HEP shot setup at 5:25 AM.

Operations established store 7074 at 6:50 AM with an average initial luminosity of 198.53E30.

Operations established beam to MTest after the store began to collide.

The Plans for Friday and the Weekend

The plans for today and the weekend are to stack, stash, and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 05/22/2009    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory